OpenStudy (anonymous):

Haii I need Help please.! <3 /.\ Evaluate the following expression using the values given: Find 4x - y - 2z if x = -2, y = 3, and z = -2 Thank You <3

OpenStudy (anonymous):

the answer is -7

OpenStudy (abmon98):

4(-2)-3-2(-2)=-7

OpenStudy (anonymous):

Thank You Dear.!! <3

jimthompson5910 (jim_thompson5910):

4x - y - 2z 4(-2) - y - 2z ... replace x with -2 (since x = -2) 4(-2) - 3 - 2z ... replace y with 3 (since y = 3) 4(-2) - 3 - 2(-2) ... replace z with -2 (since z = -2) Now evaluate and simplify 4(-2) - 3 - 2(-2) -8 - 3 + 4 -11 + 4 -7 This means 4x - y - 2z = -7 when x = -2, y = 3, and z = -2

OpenStudy (anonymous):

Can i ask another real quick.?

OpenStudy (anonymous):

i got that by putting it into my calaculator

OpenStudy (anonymous):

x over 8 - 5 = -1?

OpenStudy (abmon98):

x=32

OpenStudy (abmon98):

x/8=5-1, 8*5=32

OpenStudy (anonymous):

well you would go head and do 8-5 to get x/3=-1 then multiply by 3 on each side to get x=-3

OpenStudy (anonymous):

thank you.!!! (: im hating this algebra /.\

OpenStudy (anonymous):

id rather do this that history:))

OpenStudy (anonymous):

trade.? lol

OpenStudy (anonymous):

haha sure ill answer your algebra questions you answer my history lol

OpenStudy (anonymous):

any more questions

OpenStudy (anonymous):

Kimberley wants to plot a graph to show the price of different numbers of skateboards. She has to plot the following points: (1, 49.99), (2, 99.98), (3, 149.97), (4, 199.96) Which of the following best describes how the x-axis should be labeled? From 0 to 6 at intervals of 3 From 0 to 8 at intervals of 4 From 0 to 3 at intervals of 1 From 0 to 5 at intervals of 1

OpenStudy (anonymous):

If you need any help with history . messsage me (:

OpenStudy (anonymous):

i think it is from 0 to 5 at intervals of 1 but im not for sure